OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (–2, 0) and (2, 8)?

OpenStudy (anonymous):

First you need the slope: (y2-y1)/(x2-x1) => (8-0)/2-(-2)) = 8/4 = 2

OpenStudy (anonymous):

now we have the slope, the equation is: y = mx + b where m is the slope and b is the intercept using the slope and a point we have: 0 = 2(-2) + b 0 = -4 + b 4 = b so the equation is y = 2x + 4

OpenStudy (anonymous):

y = 2(x + 2) y = –2(x + 2) y = 2(x – 2) y = –2(x – 2)

OpenStudy (anonymous):

those are the answers given

OpenStudy (anonymous):

that's the same as y=2(x+2) (just factored the 2 out)

OpenStudy (anonymous):

ok i got it lol sorry had a dumb moment thanks tbates